OpenStudy (anonymous):

Find the minimum y-value on the graph y=f(x) f(x)= x^2+4x-5

OpenStudy (anonymous):

if you want to know how to do it by hand... you find the vertex. first step is to find the axis of symmetry, which is -b/2a. once you find that, you know the x coordinate of the vertex (which is the lowest point in the graph, since it is a positive quadratic), so all you have to do is plug in your x to find the y coordinate.

OpenStudy (anonymous):

Well when I do it by how they are showing me what I am getting and what they are getting is two COMPLETELY different things.. extremely frustrating, I used the -6/2a

OpenStudy (anonymous):

well, you know that its not negative 6 over two a, its negative b over two a. in this case, a = 1 b = 4

OpenStudy (anonymous):

*-b/2a sorry

OpenStudy (anonymous):

then f(-2) followed by 2^2+4*(-2)-5 correct therefore it gives you -1 ?

OpenStudy (anonymous):

not negative one. it should be negative 9. try your arithmetic again.

OpenStudy (anonymous):

and yet it says your answer is incorrect as well.

OpenStudy (anonymous):

maybe you typed the equation wrong? the absolute lowest y value for the equation x^2+4x-5 is -9. here's it's graph. https://www.google.com/search?q=x^2%2B4x-5&ie=utf-8&oe=utf-8&aq=t&rls=org.mozilla:en-US:official&client=firefox-a

OpenStudy (anonymous):

find \(x=-\large{b\over2a}\) given that the equation is in \(f(x)=ax^2+bx+c\) form.\[f(x)=x^2+4x-5\]\[x=-{4\over2(1)}\]\[x={-{4\over2}}\]\[x=-2\]now plug in this \(x\) value into the equation and find \(f(x)\):\[f(-2)=(-2)^2+4(-2)-5\]\[f(-2)=4+(-8)-5\]\[f(-2)=4-8-5\]\[f(-2)=-4-5\]\[f(-2)=-9\] and thus the minimum value of \(y\) is \(-9\)